OpenStudy (anonymous):

What is the 17th term of the arithmetic sequence 18, 22, 26, … ?

OpenStudy (anonymous):

a) 82 b)78 c)68 d)74

OpenStudy (anonymous):

a

OpenStudy (anonymous):

so the nth term of this(a_n) is given by a_n=a1+(n-1)d in this case the first term is 18 and the difference between two consecutive numbers on the sequence in 4 so the 17th term is a_17=18+(17-1)4=18+64=82